bob is 1/4 as old as mary in four years from now bob will be 3/4 of marys age
Answer by stanbon(75887)

bob is 1/4 as old as mary in four years from now bob will be 3/4 of marys age
-----
Equations:
b = (1/4)m
(b+4) = (3/4)(m+4)
--------
Substitute for "b" and solve for "m":
((1/4)m + 4 = (3/4)m + 3
(1/2)m = 1
m = 2 (Mary's age now)
-----
bob = (1/4)m = 6 months (Bob's age now)
==========================================
Cheers,
Stan H.
=================